# A reanalysis of the FDA’s benefit-risk assessment of Moderna’s mRNA-1273 COVID vaccine stratified not only based on age and sex but also on prior-infection and comorbidity status

### v3

@@ 2.3. The FDA’s mRNA-1273 Risks Assessment
The FDA relied on its Biologics Effectiveness and Safety (BEST) System for the VAM/P case rates appearing in Table 1 above and used the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of VAM/P cases requiring hospitalization [11, Section 2.3.2].
@@
VAM/P-case rates vary considerably depending on the case definition employed and the extent of stratification of risk, based on age, sex, vaccine manufacturer and dose, with most cases occurring postdose 2 [9]. For its mRNA-1273 assessment, the FDA four BEST-System data partners reported, with wide confidence intervals, the following myo/pericarditis incidence rates for 18–25-year-old males during the 7-day interval following dose 2: approximately 35, 70, 146, and 256 cases per million second doses [40, p, 12]. The FDA’s BEST System uses health-claims data not validated by a complete review of patients’ medical charts [11, Section 4], and has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [10, Section 4]. We rely on four sources of data to estimate the VAM/P rate for dose 2 among 18–25-year-old males: (i) Buchan et al. [41] who apply the Brighton Collaboration’s case definition, (ii) Patone et al. [42] who consider only hospitalized cases of myocarditis, (iii) Sharff et al. [43] who apply the CDC’s case definition [44], and (iv) the FDA (with VSD-based adjustments).
The FDA underestimated mRNA-1273 VAM/P risk:
@@
$\bullet$ The FDA omitted dose-one VAM/P events. For its mRNA-1273 analysis [11], the FDA computed the evaluation-period benefits of two doses of mRNA-1273 without accounting for the VAM/P risk associated with dose 1, potentially biasing the FDA’s analysis in favor of vaccination. In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [10, Section 2.3.1]. Hence, for BNT162b2, the FDA did account for 1st-dose VAM/P risk but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination, potentially biasing the analysis against vaccination. Because adverse events of vaccination other than VAM/P were disregarded by the FDA and mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit-risk analysis assessing benefits only postdose 2, the choice to include 1st-dose VAM/P risk is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in the benefit-risk analysis of its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 2.4. Thus, we include 1st-dose VAM/P risk for mRNA-1273 in our main analysis. In Supplement 2, we conduct a sensitivity analysis showing that accounting for dose 1 benefits does not substantially change our benefit-risk conclusions.
@@
$\bullet$ The FDA disregarded evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2. In a “Review Memorandum Addendum” [45], dated 11/19/21, the FDA describes multiple sources of evidence that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2, including reports from regulatory agencies of five countries, data from Ontario later analyzed by Buchan et al. [41], and data from the CDC: “[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….”
@@
The “most recent data” from the VSD to which the FDA refers was presented at the CDC’s October 2021 ACIP meeting [46]. Adjusted-rate ratios for 18–39-year-old males from the table on page 27 of [46] suggest the 2nd-dose VAM/P rate for mRNA-1273 is 2.26 times that of BNT162b2. Data from other sources suggests higher rate ratios e.g,, a 2nd-dose rate ratio of 5.1 was reported by Buchan et al. for males 18–24 [41, p. 5] and of 5.2 was reported by Husby et al. for males 12–39 [47, paragraph after Figure 3]. A study by Patone et al. of myocarditis cases requiring hospitalization within a 28-day risk interval after a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and 12 per million 1st doses, with corresponding BNT162b2 rates being 12 per million 2nd doses and 3 per million 1st doses [42, p. 4].
@@
Based on “preliminary findings” [45, p. 2], having “large uncertainty” owing to “small numbers of observed events” [40, p. 16], as well as omitting dose-1 cases, the FDA suggests that for 18–25-year-old males VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [11, Table 4] vs. 131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [10, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations. Throughout this paper, we assume that VAM/P rates for males 18–24 well approximate those for males 18–25.
@@
Using the rate ratios and BNT162b2 VAM/P-risk assessment discussed above, we compute in Section 6.1 of Supplement 1 a VAM/P hospitalization rate of 224 per million 2nd doses of mRNA-1273 for males 18–25 years old.
@@
The FDA’s BEST System also provided evidence that VAM/P risk for mRNA-1273 might be over twice that for BNT162b2. An FDA briefing document for a meeting held 10/14/2021 indicates that one of the agency’s BEST datapartners reported a postvaccination myo/pericarditis incidence rate of 283.7 (95% CI 145.2-573.5) cases per million full mRNA-1273 vaccinations (7 day risk window after each dose) [48, Section 2.3.3]. Applying a background rate of 4 cases per million per week for 18–25-year-old males [Supplement 1, Section 6.3] to each dose, we find the VAM/P case rate corresponding to 283.7 is 275.7 per million full vaccinations. Assuming a case-hospitalization rate of 100%, as in FDA’s BNT162b assessment, we obtain 275.7 VAM/P hospitalizations per million full mRNA-1273 vaccinations vs. 131 per million full BNT162b vaccinations.
$\bullet$ The FDA’s meta-analysis of its BEST-System data [11, Section 2.3.1] yielded VAM/P-rate estimates substantially lower than those of major international studies by Buchan et al. [41] and Patone et al. [42], as well as a U.S. study by Sharff et al. [43] [^5] The conclusion of [43] includes “We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm, which depends on select hospital discharge diagnosis codes. The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees.” . In Table 3 below, we compare the FDA’s projected mRNA-1273 VAM/P case and hospitalization rates to those of the other sources discussed above. (We have excluded VAM/P-rate sources, such as [49], depending on strictly passive surveillance—see Section 6.8 of Supplement 1.)
@@
a For discussions of the data in columns 3–6, see Section 6 of Supplement 1; in particular, data in column 4 are estimates based on findings from [42] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [41]. To obtain the column-5 VAM/P case rate of 525 per million we applied the background myo/pericarditis rate suggested by [41].
b Buchan et al. applied the Brighton Collaboration’s case definition for myo/pericarditisx, which reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [41, p. 5].
@@
d For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [10, Section 2.3.2.2]. For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [11, Section 2.3.2.2]. We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate: 6 of the 7 cases were hospitalized). For column 3, Buchan et al. report a case-hospitalization rate postdose 2 of an mRNA vaccine (all ages, males & females) of 79.2% ($\approx 164/207\times 100$%) [41, Table 1]; we expect the rate to be higher for the age range 18–24, but, conservatively, choose a hospitalization rate of 80%.
@@ 2.4. A Seventh Benefit Risk Scenario for mRNA-1273
The FDA’s BLA Clinical Review Memorandum [51] for mRNA-1273, completed 1/28/2022, describes a benefit-risk scenario not matching any scenario of the FDA’s published benefit-risk assessment of mRNA-1273 [11]. The projections of the BLA-Memo Scenario [51, Section 4.7] are described in the initial row of data of Table 4 below and compared to the FDA’s scenarios of [11, Table 2] in particular to Scenario 1, the most likely scenario. Differences include the BLA-Memo Scenario’s accounting for 1st-dose VAM/P risk and projecting 1 million full vaccinations would cause 47 ICU admissions (vs. 0 admissions). [^6] Subsequent analyses provide further evidence that the FDA may underestimate VAM/P severity in [11]; e.g., a meta-analysis [52] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis and a 23.2% ICU-admission rate.
@@
* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Projections for the FDA’a Scenarios 1–6 are from [11, Table 2]. BLA-Memo-scenario projections are from [51, Section 4.7].”
Consistent with the BLA Memo Scenario, the FDA’s 1/31/22 press release [12] announcing the full approval of the mRNA-1273 vaccine indicates some mRNA-1273 VAM/P cases do require ICU care. Both the BLA Memo [51, Section 4.7] and the FDA’s mRNA-1273 benefit-risk assessment [11, Section 2.3.2.2] attribute their contradictory ICU risk estimates to VSD data. The BLA Memo lacks a specific reference, while the FDA in [11] cites a source of VSD data [53] with collection cutoff point 8/21/2021, over four months before completion of the FDA’s mRNA-1273 assessment.
The VSD data from [53] cited by the FDA [11] is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 [53, p. 16]. A comparison of Buchan et al.’s myo/pericarditis data from Ontario [41, Table 4, collected 6/1/21– 9/4/21] to VSD data through 10/9/21 [46, pp. 8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million second doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system [Supplement 2, Appendix 3].
3. Reanalysis: Methods
Recall that for its case and hospitalization endpoints, the FDA’s benefit-risk model has inputs $C$ and $H$ (the numbers of evaluation-period COVID cases and hospitalizations, respectively, in a test population of one million unvaccinated males) as well as VE and VEH (the fractions of these events prevented by full vaccination). These model inputs are combined to project the cases-prevented and hospitalizations-prevented benefits of vaccinating the test population, and these benefits are compared to the model’s corresponding risk inputs, the numbers of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose.
Our model, which has hospitalizations as its endpoint, is algebraic (like the FDA’s) and thus is widely accessible. Our model extends the FDA’s by not only incorporating prior-infection protection but also by being essentially equivalent to the FDA’s model if prior infection is assumed to provide no protection. Because the FDA relied on CDC data to supply projected values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly doesn’t explicitly link VE and VEH. In contrast, our model does provide these links. For example, it links infections, a key model input replacing reported cases, to hospitalizations via another key input—the infection hospitalization rate IHR. Additional input variables are described in Section 3.1 below and summarized in Table 5.
@@ 3. Reanalysis: Methods
Test population. Our model is based on a hypothetical test population of 1 million males 18–25. We assume that the test population is representative, as of 1/1/2022, of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection. We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1/1/22 – 5/31/22 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1/1/22). The difference—the projection of (a) minus that of (b)—is the hospitalizations-prevented benefit of vaccinating the test population. We call a test population consisting of infection naïve unvaccinated persons a baseline test population.

### v4

@@ 1.4. The FDA’s mRNA-1273 Risk Assessment
The FDA relied on its Biologics Effectiveness and Safety (BEST) System for the VAM/P-case rates appearing in Table 1 above and used the CDC’s Vaccine Safety Datalink (VSD) system to estimate the percentage of cases requiring hospitalization [13, Section 2.3.2].
@@
VAM/P-case rates vary considerably depending on the case definition employed and the extent of stratification of risk, based on age, sex, vaccine manufacturer, and dose, with most cases occurring postdose 2 [9]. For its mRNA-1273 assessment, the FDA’s four BEST-System data partners reported, with wide confidence intervals, the following myo/pericarditis incidence rates for 18–25-year-old males during the 7-day interval following dose 2: approximately 35, 70, 146, and 256 cases per million 2nd doses [44, p, 12]. The FDA’s BEST System uses health-claims data not validated by a complete review of patients’ medical charts [13, Section 4] and has “inherent limitations, such as small sample sizes and imperfect sensitivity of ICD-10 codes to identify these rare outcomes” [12, Section 4]. We rely on four sources of data to estimate the VAM/P rate for dose 2 among 18–25-year-old males: (i) Buchan et al. [14], who apply the Brighton Collaboration’s case definition, (ii) Patone et al. [15], who consider only hospitalized cases of myocarditis, (iii) Sharff et al. [16], who apply the CDC’s case definition [45], and (iv) the FDA (with VSD-based adjustments).
How our mRNA-1273 VAM/P-risk assessment differs from the FDA’s:
@@
$\bullet$ We include dose-one VAM/P events. For its mRNA-1273 analysis [13], the FDA computed the evaluation-period benefits of two doses of mRNA-1273 without accounting for the VAM/P risk associated with dose 1, potentially biasing the FDA’s analysis in favor of vaccination. In contrast, for BNT162b2-related VAM/P, the FDA added excess myo/pericarditis cases postdose 1 and postdose 2 to approximate the population-level risk of completing 1 million two-dose vaccination series [12, Section 2.3.1]. Hence, for BNT162b2, the FDA did account for 1st-dose VAM/P risk but did not assess the benefits of dose 1 alone during the interval starting with dose 1 and ending at the point of full vaccination, potentially biasing the analysis against vaccination. Because we, like the FDA, do not account for adverse events of vaccination other than VAM/P, and mRNA vaccination might have unknown long-term adverse health impacts, we believe that even in a benefit–risk analysis assessing benefits only postdose 2, the choice to include 1st-dose VAM/P risk is appropriate, being consistent with the first-do-no-harm principle, the FDA’s choice in its BNT162b2 assessment, and the FDA’s choice in the benefit–risk analysis of its BLA (Biologics License Application) Clinical Review Memorandum for mRNA-1273—see Section 1.5 below. Thus, we include the 1st-dose VAM/P risk for mRNA-1273 in our main analysis. See Section S2 of Supplement S2 for a sensitivity analysis showing that accounting for dose 1’s benefits does not substantially change our benefit–risk conclusions.
@@
$\bullet$ We incorporate evidence that VAM/P risk for mRNA-1273 was over twice that for BNT162b2. In a “Review Memorandum Addendum” [46], dated 19 November 2021, the FDA describes evidence that VAM/P risk for mRNA-1273 is significantly greater than that for BNT162b2, including regulatory-agency reports from five countries, data from Ontario later analyzed by Buchan et al. [14], and data from the CDC: “[T]he most recent data presented by the Center for Disease Control and Prevention (CDC) analyzed from the Vaccine Safety Datalink (VSD) indicates a higher rate of myocarditis with the Moderna COVID-19 Vaccine than the Pfizer-BioNTech COVID-19 Vaccine ….”
@@
The “most recent data” from the VSD to which the FDA refers was presented at the CDC’s October 2021 ACIP meeting [17]. Adjusted-rate ratios for 18–39-year-old males from the table on page 27 of [17] suggest the 2nd-dose VAM/P rate for mRNA-1273 is 2.26 times that of BNT162b2. Data from other sources suggests higher rate ratios, e.g., a 2nd-dose rate ratio of 5.1 was reported by Buchan et al. for males 18–24 [14, p. 5] and of 5.2 was reported by Husby et al. for males 12–39 [47, paragraph after Figure 3]. A study by Patone et al. of myocarditis cases requiring hospitalization within a 28-day risk interval after a vaccine dose found among males under 40 a rate of 101 excess myocarditis hospitalizations per million 2nd doses of mRNA-1273 and 12 per million 1st doses, with corresponding BNT162b2 rates being 12 per million 2nd doses and 3 per million 1st doses [15, p. 4].
@@
Based on “preliminary findings” [46, p. 2], having “large uncertainty” owing to “small numbers of observed events” [44, p. 16], as well as omitting dose-1 cases, the FDA suggests that for 18–25-year-old males, VAM/P risk for mRNA-1273 is comparable to that for BNT162b2: 128 cases per million 2nd doses of mRNA-1273 [13, Table 4] vs. 131 cases per million full vaccinations of BNT162b2 (among 18–24-year-old males) [12, Table 3], with corresponding hospitalization projections of 110 per million 2nd doses of mRNA-1273 and 131 per million full BNT162b2 vaccinations. Throughout this paper, we assume that VAM/P rates for males 18–24 well approximate those for males 18–25.
@@
Using the rate ratios and BNT162b2 VAM/P-risk assessment discussed above, we compute a VAM/P-hospitalization rate of 224 per million 2nd doses of mRNA-1273 for males 18–25 years old [Supplement S1, Section S6.1].
@@
The FDA’s BEST System also provided evidence that VAM/P risk for mRNA-1273 might be over twice that for BNT162b2. An FDA briefing document for a meeting held 14 October 2021 indicates that one of the agency’s BEST data partners reported a postvaccination myo/pericarditis incidence rate of 283.7 (95% CI 145.2–573.5) cases per million full mRNA-1273 vaccinations (7 day risk window after each dose) [48, Section 2.3.3]. Applying a background rate of four cases per million per week for 18–25-year-old males for each dose [Supplement S1, Section S6.3], we find the VAM/P-case rate corresponding to 283.7 is 275.7 per million full vaccinations. Assuming a case-hospitalization rate of 100%, as in FDA’s BNT162b assessment, we obtain 275.7 VAM/P hospitalizations per million full mRNA-1273 vaccinations vs. 131 per million full BNT162b vaccinations.
As we have indicated, we rely on four sources of data for our VAM/P-rate estimates for dose 2 of mRNA-1273: the FDA with VSD-based adjustments, as well as studies by Buchan et al. [14], Patone et al. [15], and Sharff et al. [16] (from, respectively, Ontario, England, and the U.S.). The conclusion of [16], which includes the following, is noteworthy: [^6] The preprint [16] and the later publication [4] share the same conclusion.
We identified additional valid cases of myopericarditis following an mRNA vaccination that would be missed by the VSD’s search algorithm $\ldots$. The true incidence of myopericarditis is markedly higher than the incidence reported to US advisory committees.
In Table 2 below, we compare the FDA’s projected mRNA-1273 VAM/P case and hospitalization rates to those of the other sources discussed above. We exclude VAM/P-rate sources, such as [49], depending on strictly passive surveillance—see Section S6.8 of Supplement S1.
@@
a For discussions of the data in columns 3–6, see Section S6 of Supplement S1; in particular, data in column 4 are estimates based on findings from [15] for males under 40, adjusted using an age-based myocarditis-incidence ratio as well as pericarditis data from [14]. To obtain the column-5 VAM/P case rate of 525 per million we applied the background myo/pericarditis rate suggested by [14].
b Buchan et al. applied the Brighton Collaboration’s case definition for myo/pericarditisx, which reduced the reported 417 post-mRNA-vaccination study-period cases of myo/pericarditis to 297 [14, p. 5].
@@
d For its BNT162b2 assessment, the FDA assumed all VAM/P cases were hospitalized [12, Section 2.3.2.2]. For its mRNA-1273 assessment the FDA assumed VAM/P case-hospitalization rates decrease with age, starting with 86% for males 18–25 and ending with 77% for males 36 and over [13, Section 2.3.2.2]. We apply the 86% rate in columns 4–6 (which, for column 5, matches the case hospitalization rate: 6 of the 7 cases were hospitalized). For column 3, Buchan et al. report a case-hospitalization rate postdose 2 of an mRNA vaccine (all ages, males & females) of 79.2% ($\approx 164/207\times 100$%) [14, Table 1]; we expect the rate to be higher for the age range 18–24, but, conservatively, choose a hospitalization rate of 80%.
@@ 1.5. A Seventh Benefit Risk Scenario for mRNA-1273
The FDA’s BLA Clinical Review Memorandum [51] for mRNA-1273, completed 28 January 2022, describes a benefit–risk scenario not matching any scenario of the FDA’s published benefit–risk assessment of mRNA-1273 [13]. The projections of the BLA-Memo Scenario [51, Section 4.7] are described in the initial row of data of Table 3 below and compared to the FDA’s scenarios of [13, Table 2] in particular to Scenario 1, the most likely scenario. Differences include the BLA-Memo Scenario’s accounting for 1st-dose VAM/P risk and projecting 1 million full vaccinations would cause 47 ICU admissions (vs. 0 admissions). [^7] Subsequent analyses provide further evidence that the scenarios of [13] may underestimate VAM/P severity; e.g., a meta-analysis [52] of 23 studies including 854 patients ages 12–20 years found a 92.6% case-hospitalization rate for postvaccination myo/pericarditis and a 23.2% ICU-admission rate.
@@
* For Scenarios 1–6, the FDA estimated VAM/P risk only for dose 2. Projections for the FDA’a Scenarios 1–6 are from [13, Table 2]. BLA-Memo-scenario projections are from [51, Section 4.7].
Consistent with the BLA-Memo Scenario, the FDA’s 31 January 2022 press release [18] announcing the full approval of mRNA-1273 indicates some mRNA-1273 VAM/P cases require ICU care.
To support its estimates of VAM/P-hospitalization and ICU-admission risk, the FDA in [13] cites a source of VSD data [53] with a collection-cutoff point of 21 August 2021, over four months before completion of its mRNA-1273 assessment. The VSD data from [53] cited by the FDA [13] is scant, reporting only 21 postvaccination myo/pericarditis cases for those 18–29 [53, p. 16]. A comparison of Buchan et al.’s myo/pericarditis data from Ontario [14, Table 4, collected 1 June 2021–4 September 2021] to VSD data through 19 October 2021 [17, pp. 8, 16] shows that, among males and females ages 18–39, the number of myo/pericarditis cases per million 2nd doses of mRNA-1273 detected by Public Health Ontario is 3.76 times the number detected by the CDC’s VSD system [Supplement S2, Appendix S3].
2. Methods
Recall that for its cases and hospitalizations endpoints, the FDA’s benefit–risk model has inputs $C$ and $H$ (the numbers of anticipated evaluation-period COVID cases and hospitalizations, respectively, in a test population of one million unvaccinated males) as well as $VE$ and $VEH$ (the fractions of these events prevented by full vaccination). These model inputs are combined to project the cases-prevented and hospitalizations-prevented benefits of vaccinating the test population, and these benefits are compared to the model’s corresponding risk inputs, the numbers of VAM/P cases and hospitalizations projected for the 2nd mRNA-1273 dose.
Our model, which has hospitalizations as its endpoint, is algebraic (like the FDA’s) and thus is widely accessible. Our model extends the FDA’s by not only incorporating prior-infection protection but also by being essentially equivalent to the FDA’s model if prior infection is assumed to provide no protection. Because the FDA relied on CDC data to supply projected values for $C$ and $H$, the FDA’s model does not explicitly link cases to hospitalizations and similarly does not explicitly link $VE$ and $VEH$. In contrast, our model does provide these links. For example, it links infections, a key model input replacing reported cases, to hospitalizations via another key input—the infection-hospitalization rate IHR. Additional input variables are described below.
@@ 2. Methods
Test population. Our model is based on a hypothetical test population of 1 million males aged 18–25. We assume that the test population is representative, as of 1 January 2022 of all males in this age group in terms of: distribution of ages within the 18–25 range, comorbidities (e.g., obesity, cardiovascular disease, diabetes), and the percentage having had a prior COVID-19 infection. We project the number of hospitalizations for Omicron infection that will occur in this test population over the evaluation period 1 January 2022–31 May 2022 (a) assuming all members remain unvaccinated throughout the period, and (b) assuming all members of the population are vaccinated throughout the period (all fully vaccinated, 14 days past dose 2, on 1 January 2022). The difference—the projection of (a) minus that of (b)—is the hospitalizations-prevented benefit of vaccinating the test population. We call a test population consisting of infection-naïve unvaccinated persons a baseline test population.